A patient has been prescribed chlorpromazine (Thorazine) 35 mg IM STAT. The medication is available in vials of 50 mg/2 mL. How many mL should be administered to the patient? ________ mL(round to the nearest 10th)

240 mg are ordered. Medication is available in 2 grains/1 tablet. How many

tablets should be given?_____tablets 1, 2,3,or 4 tablets?
Answer #1

Ans) Order: Chlorpromazine (Thorazine) 35 mg IM STAT.

The medication is available in vials of 50 mg/2 mL.

50 mg = 2 ml

35 mg = x

x= 35 × 2/ 50 = 1.4 ml

240 mg are ordered. Medication is available in 2 grains/1 tablet.

2 grains = 129.598 milligrams ~ 130 mg

130 mg = 1 tablet

240 mg = x

x = 240/130 = 1.84 ~ 2 tablets.